OpenStudy (anonymous):

Let y = 45,000(0.97)x represent the purchasing power of $45,000 with an inflation rate of 3% per year, where x represents the number of years. What is the purchasing power when x = 0?

OpenStudy (anonymous):

\[\huge y = 45000^{0.97x}\]what do you get for x = 0, what is y?

OpenStudy (anonymous):

read the beginning

OpenStudy (anonymous):

yea y is the purchasing power when you substitute x = 0, what do you get for y?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

\[0.97 \times 0 = 0\] right?

OpenStudy (anonymous):

these are the answer choices $0 $45,000 $90,000 $43,650

OpenStudy (anonymous):

nvm what i said. i put it in wrong form \[y = 45000(0.97)^x\]\[y = 45000(0.97)^0\]\[y=45000\]

OpenStudy (anonymous):

thanks